### v3

# REVS: Unlearning Sensitive Information in Language Models via Rank Editing in the Vocabulary Space

Abstract
Language models (LMs) risk inadvertently memorizing and divulging sensitive or personally identifiable information (PII) seen in training data, causing privacy concerns. Current approaches to address this issue involve costly dataset scrubbing, or model filtering through unlearning and model editing, which can be bypassed through extraction attacks. We propose REVS, a novel non-gradient-based method for unlearning sensitive information from LMs. REVS identifies and modifies a small subset of neurons relevant for constituent tokens that form sensitive information. To adequately evaluate our method on truly sensitive information, we curate three datasets: email and URL datasets naturally memorized by the models, and a synthetic social security number dataset that we tune the models to memorize. Compared to other methods, REVS demonstrates superior performance in unlearning sensitive information and robustness to extraction attacks, while retaining underlying model integrity.
REVS: Unlearning Sensitive Information in Language Models via Rank Editing in the Vocabulary Space
Tomer Ashuach    Martin Tutek    Yonatan Belinkov Technion – Israel Institute of Technology {tomerashuach,martin.tutek,belinkov}@campus.technion.ac.il
1 Introduction
Figure 1: Overview of REVS: (1) The original model memorizes a sensitive email address and (2) reproduces it exactly given a related prompt. (3) After applying REVS, the target email token(s) are demoted to a lower rank $R$ in the model’s output, preventing the model from generating the unlearned email.
Language models (LMs) exhibit a concerning tendency to memorize information from their training data (Petroni et al., 2019; Carlini et al., 2021; Chang et al., 2023). While factual recall is a desirable property when dealing with general knowledge, memorization and regurgitation of sensitive private information, such as personal and contact details, is a security concern regulated by laws like the general data protection regulation (GDPR; European Union, 2016). To ensure such information does not get inadvertently leaked, it is paramount to develop techniques that detect and erase sensitive information from LMs or their training data.
Current approaches for handling sensitive information in LMs fall into two groups. Exact unlearning approaches tackle the problem from the data perspective, either erasing information from datasets (Kandpal et al., 2022; Lee et al., 2022) or applying differential privacy (Abadi et al., 2016; Hoory et al., 2021; Li et al., 2022; Ramaswamy et al., 2020) to training data. Such approaches are costly and time-consuming, as each iteration of scrubbing requires retraining the entire model. Moreover, they reduce but do not entirely prevent the risk of information leakage (Carlini et al., 2019, 2023a).
Machine unlearning approaches (Liu et al., 2024) discourage models from generating sensitive information through in-context unlearning (ICL; Madaan et al., 2022; Pawelczyk et al., 2024; Zheng et al., 2023) or gradient ascent (Jang et al., 2023; Yao et al., 2023; Yu et al., 2023). These approaches do not ensure erasure of sensitive information from model parameters, rendering them vulnerable to extraction attacks (Li et al., 2023). Model editing, a variant of localization-based unlearning, localizes and edits a subset of parameters to erase sensitive information (Cao et al., 2021; Meng et al., 2022, 2023). In contrast to ICL and optimization-based methods, which prevent models from generating sensitive content but do not fundamentally erase the underlying knowledge, editing methods hold greater potential to resist extraction attacks (Carlini et al., 2021).
We introduce Rank Editing in the Vocabulary Space (REVS), a novel non-gradient-based model-editing approach that enables robust unlearning of sensitive information from LMs while maintaining model performance and offering strong robustness against extraction attacks. Adopting the view that transformer MLP layers construct predictions by promoting specific tokens in the output vocabulary space (Geva et al., 2022), REVS locates the layers and particular subsets of neurons that promote the tokens corresponding to the targeted sensitive information. By modifying these neurons with respect to the relevant tokens, REVS surgically removes the model’s encoded tendency to generate that sensitive data while preserving its broader knowledge and remaining robust to extraction attacks. See Figure 1 for an illustration.
We aim to prevent models from generating specific token sequences rather than erasing broader conceptual knowledge evaluated in datasets such as WMDP (Li et al., 2024) and TOFU (Maini et al., 2024). To this end, we curate three benchmark datasets containing sensitive information: Emails and URLs, containing real addresses inadvertently memorized by the model during pretraining, and SSN, a synthetic dataset where we instilled social security numbers into LMs through finetuning. Unlike prior work (Patil et al., 2024), which evaluated editing methods on non-sensitive data, our benchmarks contain actual private information, enabling rigorous assessment of unlearning efficacy and robustness against extraction attacks. We conduct experiments on Llama-3-8B (Dubey et al., 2024) and GPT-J-6B (Wang and Komatsuzaki, 2021), which we find to naturally memorize sensitive information from their training data (The Pile; Gao et al., 2020). Our experiments demonstrate that REVS outperforms six strong baselines in unlearning sensitive information while maintaining model performance and robustness against extraction attacks. Our main contributions are:
  • We introduce REVS, a novel method for unlearning sensitive information from LM.
  • We curate three datasets containing sensitive information, two of which are naturally memorized.
  • We demonstrate REVS’s superior performance in unlearning efficacy and robustness against extraction attacks while preserving the model’s general capabilities.
2 Problem Setup and Preliminaries
Consider the case where a LM is prompted with text from its training dataset that contains sensitive information, e.g., “Contact David Lewis at ”. If the model has memorized the completion of this prompt, it will generate the relevant email address, such as “lewis.david@email.com”, illustrated in Figure 1. This scenario highlights a prevalent issue with LMs: their ability to memorize can lead to unintended divulgence of sensitive data, in particular PII. To address this issue, dataset scrubbing approaches tackle it from a data perspective, preventing sensitive information from being stored in the model during training. Machine unlearning approaches take on a post-hoc approach, modifying either all, or just a subset of parameters that contain sensitive information. The goal of these approaches is to efficiently and effectively prevent the model from generating each sequence containing sensitive information in a way robust to extraction attacks, while preserving general capabilities of the model (Liu et al., 2024).
Background on Transformer LMs.
We briefly describe the main components in Transformer LMs; more details are found elsewhere (Vaswani et al., 2017; Black et al., 2021). Omitting some technical details, a Transformer LM is made of a sequence of blocks. Each block contains self-attention and multi-layer perceptrons (MLPs) that write and read to the residual stream (Elhage et al., 2021). The MLP is defined as $\text{MLP}(\vec{x})$ $=\text{FF}_{2}\sigma(\text{FF}_{1}\vec{x})$, and is added to the residual stream. The final layer hidden state is projected to the vocabulary space via the unembedding matrix $U$, followed by a softmax to obtain next token probabilities.
Due to the residual stream tying hidden spaces across layers, applying the unembedding matrix to residual states at intermediate layers results in meaningful token logits (Nostalgebraist, 2020), while applying it to MLP weights reveals stored knowledge (Dar et al., 2023). In fact, the second MLP layer, $\text{FF}_{2}$, acts as a memory store (Geva et al., 2021; Meng et al., 2022), and model editing methods often target this layer to update a model’s knowledge (Meng et al., 2022, 2023). We adopt this view, and focus on information stored in $\text{FF}_{2}$ columns, henceforth neurons, for unlearning.
3 Methodology
Figure 2: Editing one neuron with REVS: (1) The neuron is projected from hidden space to vocabulary logit space. (2) The logit is adjusted to demote the target token rank to a desired lower rank $R$. (3) The adjusted logits vector is projected back to hidden space, yielding the updated neuron value.
We prompt a transformer LM with a prompt $P$ and observe its behavior when generating the next token. Let $\vec{a}=\sigma(FF_{1}\vec{x})$ be the intermediate representation of a transformer feed-forward layer, while $\vec{h}=FF_{2}\vec{a}$ is the output, for a given input $\vec{x}$. Leveraging the insight that applying the unembedding matrix $U$ to hidden states of intermediate layers produces meaningful logits, we define $r_{t}$ as the rank of a target token $t$ among other logits in the vocabulary projection of hidden state $h$:
$r_{t}\coloneq\text{rank}(t,U\vec{h})$ (1)
The higher the rank, the more likely the token is to be generated.
To unlearn sensitive information, REVS aims to decrease $r_{t}$ to a desired rank $r_{d}$ by identifying and modifying $FF_{2}$ columns that contribute most to generating the target token $t$. This unlearning process operates in two main phases: localization (Section 3.1), where we identify the layers and neurons ($FF_{2}$ columns) that contain information relevant for generating $t$, and editing (Section 3.2), where we overwrite information relevant for $t$ while not affecting knowledge used for generating other tokens. We provide a high-level overview of REVS in Algorithm 1.
1:procedure UnlearnToken($M,P,t,r_{d},n_{max}$)
2:     Input: Model $M$, prompt $P$, target token $t$,
3:      target desired rank $r_{d}$, max neurons $n_{\mathrm{max}}$
4:     for $l\in\mathrm{SelectLayers}(M,P,t)$ do $\triangleright$ Sec.3.1.1
5:         $r_{t}\leftarrow\mathrm{TokenRank}(M,P,l,t)$
6:         $i\leftarrow 0$
7:         while $r_{t}>r_{d}$ and $i<n_{\mathrm{max}}$ do
8:              $\vec{n}\leftarrow\mathrm{SelectNeuron}(M,P,l,t)$triangleright Sec.3.1.2
9:              $\vec{n}\leftarrow\mathrm{EditNeuron}(\vec{n},t)$triangleright Sec.3.2
10:              $r_{t}\leftarrow\mathrm{TokenRank}(M,P,l,t)$triangleright Eq.1
11:              $i\leftarrow i+1$
12:         end while
13:     end for
14:end procedure
Algorithm 1 REVS: Rank Editing in the Vocabulary Space
3.1 Neuron Localization
To effectively unlearn sensitive information while minimizing model disruption, we identify specific model components that strongly contribute to generating token $t$. While information in transformer models is distributed across multiple layers and parameters within them, prior work shows it can be effectively localized (Wu et al., 2023; Dai et al., 2022). Our localization process operates in two steps: identifying relevant layers, and selecting a small subset of neurons ($FF_{2}$ columns) within them that contain relevant information.
3.1.1 Layer Selection
We first identify relevant layers by measuring how strongly each layer contributes to generating the target token $t$. For each layer $l$, we compute the rank of the target token $r_{t,l}$ for $\vec{h}_{l}$. Layers where $r_{t,l}$ exceeds a predetermined threshold $r_{h}$ are selected as for editing. Henceforth, we omit layer indices for brevity, though the process is performed layer by layer.
3.1.2 Neuron Selection
We identify neurons to edit using two criteria:
1. Activation strength: We measure how strongly a neuron $\vec{n}_{j}$, corresponding to a $FF_{2}$ column $j$, is activated in response to prompt $P$ using $a_{j}$, where $\vec{a}=\sigma(FF_{1}\vec{x})$. A higher activation $a_{j}$ means that the information from the associated column is more strongly represented in the resulting hidden state $\vec{h}=FF_{2}\vec{a}$.
2. Token association: We measure a neuron’s association with token $t$ by computing the token’s rank $r_{t}$ when projecting the neuron to the vocabulary space $r_{t}=\text{rank}(t,U\vec{n}_{j})$.
We select neurons to edit by first identifying the top $k$ neurons with the highest activation values, then progressively selecting neurons with the strongest target token association until either $r_{t}<r_{h}$ or the maximum allowed number of edited neurons $n_{max}$ is reached in the layer.
This hybrid selection approach targets neurons that are both contextually significant and semantically relevant, thus outperforming alternatives such as selection based solely on activations, token associations, gradients (Dai et al., 2022), or random selection (see Section D.1).
3.2 Neuron Editing
For each selected neuron $\vec{n}_{j}$, REVS performs iterative adjustment as illustrated in Figure 2:
  1. 1.
    Project to vocabulary space: $\vec{v}=U\vec{n}_{j}$.
  2. 2.
    Set target token’s logit: $\vec{v}_{t}=l_{t}$.
  3. 3.
    Project back to neuron space: $\vec{n}_{j}^{*}=U^{\dagger}\vec{v}$, where $U^{\dagger}$ is the pseudoinverse of the unembedding matrix.
  4. 4.
    Set $FF_{2,j}=\vec{n}_{j}^{*}$.
We initialize $l_{t}$ with a small logit value corresponding to a low rank, and iteratively decrease it by multiplying with a constant factor in each iteration until $r_{t}<r_{n}$, where $r_{n}$ is the neuron’s desired rank. For a detailed description of the iterative process see Section B.2.
3.3 Target Token Selection
Sensitive information typically spans multiple tokens, but unlearning every token is unnecessary, as accurately recovering the original data requires reconstructing the full token sequence. To account for this, for each target sequence $S$, we select a subset of the $T\subseteq S$ rarest tokens, as unlearning targets for all considered methods, with $|T|=2$ in this work. See Sections E.1 and E.2 for details and ablation experiments.
4 Experimental Setup
4.1 Models
We use Llama-3-8B (Dubey et al., 2024) and GPT-J-6B (Wang and Komatsuzaki, 2021), which we find have memorized sensitive information from their training on The Pile (Gao et al., 2020).
4.2 Datasets
Organically memorized data: Identifying naturally memorized sensitive information requires two computationally intensive steps: detecting genuine sensitive information and verifying its memorization. Due to computational constraints, we use a pre-filtered subset of The Pile (Gao et al., 2020) containing 15,000 sentences.111github.com/google-research/lm-extraction-benchmark From this, we identified memorized instances in Llama-3-8B (205 email addresses, 203 URLs) and GPT-J-6B (288 email addresses). Synthetic SSN dataset: We generated 200 sentences containing social security numbers using Claude 3 Sonnet (Anthropic, 2024) and fine-tuned the base models to memorize them. See Appendix F for more details.
4.3 Baselines
Model editing.
Model editing techniques perform localized parameter updates to specific modules within LMs that encode knowledge about target concepts. From this family of methods, we use MEMIT (Meng et al., 2023), Head-Projection and Max-Entropy, the two best performing methods from Patil et al. (2024). We modify the objective of MEMIT to decrease the probability of generating specific target tokens in order to facilitate unlearning.
Optimization-based.
Optimization-based methods define a target loss term and update model parameters through gradient descent. We select three strong methods from this family as baselines: Constrained Fine-tuning (FT-L) (Zhu et al., 2020), NPO-KL (Zhang et al., 2024), and RMU (Li et al., 2024).
We elaborate on the used baselines and their hyperparameters in Appendix C.
4.4 Evaluation Metrics
We comprehensively evaluate the unlearning process through three critical dimensions: effectiveness (4.4.1), which measures how successfully sensitive information is unlearned from the model, integrity (4.4.2), which assesses the impact of unlearning on unrelated knowledge and model capabilities, and extraction resistance (4.4.3), measuring the model’s resistance to attempts at recovering the unlearned sensitive information through extraction attacks.
4.4.1 Unlearning Effectiveness
To evaluate unlearning effectiveness, simply reducing the probability of $t$ being the top prediction is insufficient, as sampling multiple times from the top-$k$ tokens may still generate $t$. In subsequent metric definitions, we use the idea of a capped rank score, a normalized rank metric that considers the top $k$ most probable tokens. Given a token $t$, let $r_{t}$ denote its rank in the model’s predicted token distribution. Then:
[…]
A higher score indicates that the sensitive token $t$ is less likely to be generated given the corresponding prompt, with a score of $1$ indicating $t$ is outside the top-$k$ most probable tokens.
Efficacy:
Since accurately recovering the sensitive information requires reconstructing the full token sequence, we define Efficacy@$k$ as the maximum Score@$k$ across the unlearned token subset $T$, capturing the difficulty of extracting the complete sequence:
[…]
Generalization: We evaluate unlearning generalization by measuring Efficacy@$k$ on the target token set $T$ using prompts that were unseen during the unlearning process but are tuned to generate the same sensitive information (see Section F.2, Table 12 for examples).
4.4.2 Model Integrity
We assess the potential side effects of unlearning methods on the model’s overall functionality through two key metrics:
Specificity
measures the unlearning method’s precision by evaluating its impact on tokens that should remain unaffected. We use a half of the dataset as a retain set, as these tokens share similar properties with the unlearned ones.
$\textrm{Specificity}=\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}[M_{\hat{\theta}}(P_{i})=S_{i}]$ (4)
where $P_{i}$ is a prompt and $S_{i}$ is the original memorized token sequence, such that before unlearning, $M_{\theta}(P_{i})=S_{i}$.
General capabilities:
We evaluate the models performance on the MMLU (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021) benchmarks before and after unlearning to measure degradation in general knowledge capabilities using lm-eval-harness (Gao et al., 2024).
4.4.3 Extraction Resistance
Editing and unlearning can leave trace information in the model making the erased information recoverable. Extraction resistance metrics quantify the susceptibility of erased sensitive tokens to be identified through various attacks. Given a candidate set of tokens $C$ identified by an adversary, we define the resistance to each attack as the minimum of the average efficacy score (Eq. 3) across all model layers $L$:
[…]
An attack is successful if the target is extracted from any layer. Compared to Patil et al. (2024), the extraction attacks used in this work are stricter and more comprehensive, considering more candidate tokens across all layers.
We consider the following adversarial attacks:
Logit-Lens Attack (LLA): Considers the top-$k$ and bottom-$k$ tokens in the vocabulary logit vector $\vec{v}$ as candidates, obtained by projecting each layer’s residual hidden state to the vocabulary logits:
$C_{\textrm{LLA}}=\bigcup_{\ell\in L}\textrm{top-$k$}(\vec{v}^{\ell})\cup\textrm{bottom-$k$}(\vec{v}^{\ell})$ (6)
Delta Attack (DA): Considers the top-$k$ tokens with the largest absolute changes in the vocabulary logit vector $\vec{v}$ between consecutive layers as candidates:
$C_{\textrm{DA}}=\bigcup_{\ell\in L}\textrm{top-$k$}(|\vec{v}^{\ell+1}-\vec{v}^{\ell}|)$ (7)
Method Unlearning Score $\uparrow$ Efficacy@100 $\uparrow$ General.@100 $\uparrow$ Specificity $\uparrow$ MMLU $\uparrow$ GSM8K $\uparrow$
SSN Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $100{\pm 0.00}$ 61.0561.0561.05{}61.05 $47.83$
FT-L $\underline{36.98{\pm 11.97}}$ $\underline{63.88{\pm 9.88}}$ $\underline{50.35{\pm 10.76}}$ $24.33{\pm 9.78}$ 60.9960.9960.99{}60.99 $46.62$
MEMIT $24.72{\pm 7.21}$ $30.70{\pm 9.67}$ $23.90{\pm 7.61}$ $22.67{\pm 6.50}$ 61.0261.0261.02{}61.02 $46.17$
Max-Entropy $5.12{\pm 2.13}$ $5.17{\pm 3.00}$ $3.92{\pm 2.33}$ $1.40{\pm 0.60}$ 61.0661.0661.06{}61.06 $47.46$
Head-Projection $2.98{\pm 0.79}$ $3.08{\pm 1.23}$ $2.95{\pm 0.68}$ $4.17{\pm 2.41}$ 61.0661.0661.06{}61.06 $46.92$
RMU $16.42{\pm 9.10}$ $13.47{\pm 8.42}$ $16.67{\pm 10.41}$ $\underline{38.67{\pm 14.92}}$ 60.8360.8360.83{}60.83 $48.21$
NPO-KL $11.95{\pm 4.87}$ $38.78{\pm 18.34}$ $36.13{\pm 16.59}$ $6.33{\pm 3.68}$ 61.0161.0161.01{}61.01 $47.23$
REVS (ours) $\bm{89.58^{\dagger}{\pm 1.99}}$ $\bm{98.88{\pm 1.28}}$ $\bm{89.67{\pm 3.78}}$ $\bm{82.17{\pm 5.08}}$ 60.8760.8760.87{}60.87 $44.20$
Emails Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$- $\bm{100{\pm 0.00}}$ 62.1762.1762.17{}62.17 47.9947.9947.99{}47.99
FT-L $\underline{50.30{\pm 3.04}}$ $52.98{\pm 4.23}$- ${49.25{\pm 8.50}}$ 62.1562.1562.15{}62.15 $50.94$
MEMIT $35.43{\pm 4.30}$ $63.63{\pm 3.50}$- $24.84{\pm 4.20}$ 62.2262.2262.22{}62.22 $50.64$
Max-Entropy $31.08{\pm 3.30}$ $\bm{69.75{\pm 6.30}}$- $20.22{\pm 3.10}$ 62.1162.1162.11{}62.11 $50.64$
Head-Projection $30.80{\pm 3.90}$ $\underline{64.33{\pm 4.90}}$- $20.43{\pm 3.40}$ 62.1062.1062.10{}62.10 $50.19$
RMU $17.47{\pm 3.60}$ $15.08{\pm 5.90}$- $32.58{\pm 16.00}$ 62.1062.1062.10{}62.10 $46.39$
NPO-KL $32.75{\pm 2.70}$ $24.27{\pm 3.00}$- $\underline{50.97{\pm 2.00}}$ 62.0562.0562.05{}62.05 $48.67$
REVS (ours) $\bm{62.37^{\dagger}{\pm 2.30}}$ $59.65{\pm 3.95}$- $65.70{\pm 3.79}$ 61.7761.7761.77{}61.77 47.4647.4647.46{}47.46
URL FT-L $\underline{28.03{\pm 3.95}}$ $\underline{59.13{\pm 7.71}}$- ${18.63{\pm 3.52}}$ $62.14$ $50.72$
MEMIT $17.52{\pm 4.10}$ $34.37{\pm 10.80}$- $11.98{\pm 3.00}$ $62.14$ $49.96$
Max-Entropy $12.78{\pm 3.90}$ $32.88{\pm 7.90}$- $8.06{\pm 2.80}$ $62.19$ $49.50$
Head-Projection $11.28{\pm 3.90}$ $26.32{\pm 8.40}$- $7.30{\pm 2.70}$ $62.14$ $49.81$
RMU $13.48{\pm 6.00}$ $12.22{\pm 12.20}$- $\underline{41.83{\pm 15.30}}$ $62.02$ $49.81$
NPO-KL $17.80{\pm 6.60}$ $10.97{\pm 4.40}$- $\bm{50.87{\pm 14.10}}$ $62.13$ $49.88$
REVS (ours) $\bm{44.25^{\dagger}{\pm 5.01}}$ $\bm{78.22{\pm 6.04}}$- $30.94{\pm 4.11}$ $62.31$ 47.7647.7647.76{}47.76
Table 1: Unlearning Effectiveness and Model Integrity on Llama-3-8B for $k=100$. Best results are bold, second-best underlined. REVS achieves superior Unlearning Score values across all datasets compared to baselines, where indicates statistical significance ($p<0.05$) in Wilcoxon signed-rank test.
Perturbation Attack (PA): A white-box attack that inserts random characters to the original prompts. The candidate set $C_{\textrm{PA}}$ is obtained as in $C_{\textrm{LLA}}$, but using the perturbed prompts. Appendix G provides more details.
4.5 Implementation Details
In each experiment, we unlearn all sensitive sequences before evaluation. The Unlearning Score is the harmonic mean of efficacy, specificity, and generality (for the SSN dataset), while the Resistance Score is the harmonic mean of all extraction attacks. Hyperparameters (HPs) are optimized per method on the SSN and Email datasets to maximize Unlearning Score with $k=100$, using a fixed seed (00). In the URL dataset, we use the same HPs as the Emails dataset, to assess robustness and generalization across naturally memorized PII types. We cross-validate over 6 random splits (seeds 11116666). Each split has a forget set of $50$ targets in Emails and URLs, while in SSN, it comprises half the data. The forget set is used for efficacy, generalization (SSN), and extraction resistance, while the rest is used for specificity.
5 Results
Figure 3: Results for Extraction Resistance on Llama-3-8B for $k=100$. REVS is more robust to extraction attacks. We report the full results in tabular format in Table 2 (Section A.2).
Figure 4: Llama-3-8B; Top row: SSN dataset; Middle row: Email dataset; Bottom row: URL dataset. From left to right: (a+d+g) efficacy vs. candidates size, (b+e+h) efficacy vs. specificity trade-off, (c+f+i) Unlearning Score vs. number of targets. Mean results with confidence intervals.
As Table 1 and Figure 3 show, REVS outperforms all baselines in Unlearning Score and is on par or better in Resistance Score, for Llama-3-8B across all datasets. Results on GPT-J-6B follow the same trends (Section A.3).
Effectiveness.
REVS achieves the best efficacy and generality on SSN and URLs and strong efficacy on Emails. As shown in Figure 4, perfect efficacy is possible but compromises specificity. Notably, methods surpassing REVS in efficacy on Emails achieve a lower Unlearning Score due to poor specificity.
Model Integrity.
REVS preserves model integrity while achieving strong Unlearning Score. It attains the best specificity for SSN and Emails and remains competitive on URLs. MMLU and GSM8K scores remain stable across methods, except for a GSM8K drop on SSN, likely due to the numerical nature of the targets and REVS’s significantly higher Effectiveness. Overall, we find that unlearning does not substantially impact the model’s core knowledge. However, as global metrics may miss local perturbations, specificity is particularly important for assessing unlearning.
Extraction Resistance.
As shown in Figure 3 and Table 2, REVS achieves the highest Resistance Score across all datasets and models, except for Emails on Llama-3-8B, where it ranks a close second. Notably, Tables 2 and 4 (Section A.2) highlight a clear link between higher Effectiveness and improved Resistance Score. While strong Effectiveness often compromises specificity, REVS maintains both, achieving the best or second-best Resistance Score and demonstrating balanced robust unlearning.
Qualitative examples.
Samples of generated text after unlearning are in Section B.3, showing the model generates coherent alternative text instead of memorized private emails.
5.1 Analysis
In this section we compare the robustness of REVS with FT-L, the second-best performing method, under different hyperparameters. Figure 4 shows that REVS Pareto-dominates FT-L in all but one experiment.
Robustness to candidate token size.
REVS consistently outperforms FT-L as we increase the size of the candidate token set $C$ (Figures 4, 4 and 4). While for the Emails and URLs datasets both methods exhibit similar trends, REVS maintains superior performance for each size of $C$. For the SSN dataset, REVS demonstrates near-perfect efficacy across all candidate token set sizes.
Efficacy/Specificity trade-off.
While both REVS and FT-L are affected by this trade-off (Figures 4, 4 and 4), REVS fares better, robustly maintaining high specificity when increasing edit strength. Notably, on SSN, there is almost no trade-off, with REVS achieving near-perfect specificity and efficacy.
Impact of the number of edits.
On SSN (Figure 4) REVS maintains a stable Unlearning Score across edit counts, while FT-L performs well only at its tuned hyperparameters, degrading elsewhere. Organic data (Figures 4 and 4) show a downward trend in Unlearning Score for both methods as edits increase; however, REVS consistently outperforms FT-L.
6 Related Work
Memorization of training data in neural models.
Foundation models of language (Carlini et al., 2019, 2021; Lee et al., 2022; Zhang et al., 2023), vision (Carlini et al., 2023b), and code (Ippolito et al., 2023) are able to reproduce sensitive personal information such as credit card numbers, email addresses, and API keys from training data. Carlini et al. (2023a) demonstrated that the GPT-J model, trained on the Pile, memorizes at least $1\%$ of its training data. Memorization is shown to increase with model scale (Carlini et al., 2023a) and specialized prompts can bypass guardrails to retrieve even more information than previously estimated (Carlini et al., 2023b; Ippolito et al., 2023; Rao et al., 2024), highlighting the need for erasing PII from models.
Defenses against leakage of sensitive information.
Privacy concerns have given way to development of methods preventing regurgitation of PII. Differential privacy (DP; Dwork et al., 2014; Ramaswamy et al., 2020; Hoory et al., 2021; Li et al., 2022) and dataset scrubbing through deduplication (Lee et al., 2022; Kandpal et al., 2022; Carlini et al., 2023a) prevent sensitive information from ever being stored in the model. However, using DP-SGD frequently results in worse generative models (Anil et al., 2022), and LMs can memorize even singleton training instances (Debenedetti et al., 2024). Scrubbing approaches are also expensive, requiring retraining for each new identified PII. To alleviate high cost incurred by retraining models, post-hoc model unlearning techniques (Liu et al., 2024) such as gradient ascent (Bourtoule et al., 2021; Neel et al., 2021; Eldan and Russinovich, 2023), reinforcement learning through human feedback (RLHF; Ouyang et al., 2022), and model editing (Cao et al., 2021; Meng et al., 2022, 2023; Wu et al., 2023) were devised. Most similar to ours is the work of Wu et al. (2023), who identify and zero out neurons responsible for memorized sensitive information. REVS diverges in key aspects, significantly outperforming naive zeroing by modifying selected neurons to demote relevant tokens in the vocabulary space, thus removing the tendency to generate sensitive information while preserving model integrity (see Section E.2).
Extraction attacks.
Membership inference attacks (Dwork et al., 2006; Shokri et al., 2017) try to verify whether a specific datum was in a model’s training set, while attribute inference attacks (Fredrikson et al., 2014) retrieve information on attributes of training instances. Data extraction attacks aim to retrieve full training instances. Generative LMs are susceptible to such attacks given their propensity to memorize training data (Carlini et al., 2019, 2023b). Patil et al. (2024) show that despite efforts to scrub sensitive information from models, different attacks can still locate such information in LMs. Black-box attacks retrieve information without access to model internals (Henderson et al., 2018; Lukas et al., 2023; Krishna et al., 2024), while white-box attacks use model parameters and probing (Nostalgebraist, 2020; Geva et al., 2022) to even detect information erased by model editing techniques (Patil et al., 2024). REVS is specifically designed with such attacks in mind, removing sensitive information effectively while maintaining robustness against potential attacks.
7 Conclusion
We introduce REVS, a novel unlearning method which demonstrates superior performance to state-of-the-art baselines in effectively unlearning sensitive information from LMs while preserving general capabilities and remaining robust to extraction attacks. We curate three datasets containing sensitive information, two naturally memorized by the models. Through detailed analyses we show REVS’ robustness to varying candidate token set sizes, number of unlearned targets and the trade-off between efficacy and specificity.
Limitations
While REVS shows promising results, several areas require further work: (1) scalability of unlearning to larger amounts of target information, (2) generalization to unlearning other types of PII, particularly numeric, while preserving model integrity, (3) evaluation against more fine-grained extraction attacks, including targeted analysis of MLP and Attention modules, as well as probe-based assessments of residual information, (4) evaluation on datasets in languages other than English.
Acknowledgements
This research has been supported by an AI Alignment grant from Open Philanthropy, the Israel Science Foundation (grant No. 448/20), and an Azrieli Foundation Early Career Faculty Fellowship. We would also like to express our gratitude to the Technion computer science NLP group for their invaluable consultation and assistance in improving this work.
Appendix A Results
A.1 Perturbation Attack Analysis.
As shown in Tables 2 and 4, the Perturbation Attack (PA) score is relatively low before unlearning, suggesting that the target tokens are already highly ranked in the model’s output given the perturbed prompt. After unlearning, the PA score increases significantly. Notably, for REVS, the PA score doubles post-unlearning, indicating its effectiveness against this attack. Interestingly, on Llama-3-8B for the URL dataset, some methods (e.g., RMU and NPO) exhibit a lower PA score than the unedited model, suggesting that internal memorization persists and becomes more pronounced when the original prompt is used for unlearning with these methods.
A.2 Results on Llama-3-8B
Full Results of Extraction Resistance.
We provide the complete table of results for extraction resistance on Llama-3-8B along with the standard deviations, complementing the visual summary presented in the main paper in Table 2.
Method Resistance Score $\uparrow$ Logit Lens@100 $\uparrow$ Delta@100 $\uparrow$ Perturb@100 $\uparrow$
SSN Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $96.80{\pm 0.59}$ $92.13{\pm 5.15}$
FT-L $\underline{82.77{\pm 5.73}}$ $\underline{63.88{\pm 9.88}}$ $98.08{\pm 0.57}$ $\underline{98.22{\pm 1.52}}$
MEMIT $55.20{\pm 11.98}$ $30.70{\pm 9.67}$ $97.90{\pm 0.47}$ $98.18{\pm 1.86}$
Max-Entropy $13.62{\pm 7.33}$ $5.17{\pm 3.00}$ $98.17{\pm 0.14}$ $95.30{\pm 3.20}$
Head-Projection $8.62{\pm 3.26}$ $3.08{\pm 1.23}$ $97.88{\pm 0.20}$ $94.35{\pm 4.52}$
RMU $29.50{\pm 16.43}$ $13.47{\pm 8.42}$ $98.00{\pm 0.42}$ $96.90{\pm 3.74}$
NPO-KL $61.63{\pm 15.57}$ $38.78{\pm 18.34}$ $\underline{98.47{\pm 0.24}}$ $95.08{\pm 3.73}$
REVS (ours) $\bm{99.27{\pm 0.35}}$ $\bm{98.88{\pm 1.28}}$ $\bm{98.92{\pm 0.43}}$ $\bm{100.00{\pm 0.00}}$
Emails Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $91.15{\pm 0.17}$ $36.55{\pm 4.46}$
FT-L $64.77{\pm 3.93}$ $52.98{\pm 4.23}$ $\underline{92.27{\pm 0.58}}$ $60.65{\pm 6.20}$
MEMIT $70.55{\pm 4.00}$ $63.63{\pm 3.50}$ $92.12{\pm 1.10}$ $\underline{62.95{\pm 6.20}}$
Max-Entropy $\bm{72.77{\pm 3.90}}$ $\bm{69.75{\pm 6.30}}$ $\bm{92.47{\pm 1.00}}$ $62.47{\pm 3.60}$
Head-Projection $70.28{\pm 3.70}$ $\underline{64.33{\pm 4.90}}$ $92.10{\pm 0.60}$ $61.57{\pm 4.10}$
RMU $29.82{\pm 8.40}$ $15.08{\pm 5.90}$ $91.05{\pm 0.90}$ $48.23{\pm 4.00}$
NPO-KL $42.05{\pm 3.60}$ $24.27{\pm 3.00}$ $91.17{\pm 0.40}$ $52.90{\pm 4.10}$
REVS (ours) $\underline{71.95{\pm 2.25}}$ $59.65{\pm 3.97}$ $91.85{\pm 0.62}$ $\bm{71.47{\pm 1.65}}$
URLs Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $91.00{\pm 0.00}$ $36.00{\pm 0.00}$
FT-L $\underline{56.70{\pm 10.16}}$ $\underline{58.80{\pm 7.37}}$ $\bm{95.58{\pm 0.44}}$ $\underline{41.18{\pm 13.90}}$
MEMIT $45.30{\pm 10.30}$ $34.37{\pm 10.80}$ $92.90{\pm 0.90}$ $39.35{\pm 9.60}$
Max-Entropy $44.93{\pm 7.90}$ $32.88{\pm 7.90}$ $\underline{94.92{\pm 0.50}}$ $39.50{\pm 6.80}$
Head-Projection $37.83{\pm 9.50}$ $26.32{\pm 8.40}$ $93.63{\pm 1.80}$ $33.63{\pm 7.50}$
RMU $16.88{\pm 15.10}$ $12.22{\pm 12.20}$ $92.48{\pm 1.30}$ $13.88{\pm 13.10}$
NPO-KL $16.77{\pm 6.80}$ $10.97{\pm 4.40}$ $91.08{\pm 0.40}$ $13.50{\pm 6.30}$
REVS (ours) $\bm{82.80{\pm 3.94}}$ $\bm{78.22{\pm 6.04}}$ $91.97{\pm 0.15}$ $\bm{79.75{\pm 4.96}}$
Table 2: Results for Extraction Resistance on Llama-3-8B for $k=100$. Best results are in bold, second best underlined. REVS is more robust to extraction attacks.
A.3 Results on GPT-J-6B
Method Unlearning Score $\uparrow$ Efficacy@100 $\uparrow$ General.@100 $\uparrow$ Specificity $\uparrow$
SSN Unedited $0{\pm 0.0}$ $0{\pm 0.0}$ $0{\pm 0.0}$ $100{\pm 0}$
FT-L $40.00{\pm 13.94}$ $72.05{\pm 6.34}$ $38.68{\pm 7.99}$ $39.33{\pm 22.28}$
MEMIT $\underline{78.07{\pm 2.2}}$ $\underline{98.5{\pm 2.33}}$ $\underline{61.15{\pm 3.25}}$ $84.17{\pm 3.07}$
Max-Entropy $19.88{\pm 6.57}$ $32.90{\pm 9.61}$ $13.22{\pm 6.24}$ $27.67{\pm 5.06}$
Head-Projection $43.02{\pm 11.73}$ $80.22{\pm 2.28}$ $52.68{\pm 4.96}$ $29.33{\pm 12.57}$
RMU $12.23{\pm 4.18}$ $19.90{\pm 6.08}$ $17.67{\pm 8.88}$ $22.17{\pm 23.56}$
NPO-KL $37.57{\pm 4.55}$ $52.95{\pm 10.80}$ $27.53{\pm 5.62}$ $47.17{\pm 12.35}$
REVS (ours) $81.45{\pm 3.56}$ $99.95{\pm 0.07}$ $80.17{\pm 3.22}$ $\underline{70.33{\pm 7.84}}$
Emails Unedited $0{\pm 0.0}$ $0{\pm 0.0}$- $100{\pm 0}$
FT-L $25.67{\pm 3.60}$ $21.52{\pm 2.29}$- $32.41{\pm 7.83}$
MEMIT $70.05{\pm 1.16}$ $\underline{88.23{\pm 1.64}}$- $58.1{\pm 1.63}$
Max-Entropy $43.6{\pm 2.05}$ $34.6{\pm 1.93}$- $57.92{\pm 3.56}$
Head-Projection $\underline{78.83{\pm 2.10}}$ $82.55{\pm 2.63}$- $\bm{75.58{\pm 3.24}}$
RMU $15.08{\pm 5.56}$ $15.93{\pm 13.72}$- $44.79{\pm 20.51}$
NPO-KL $41.07{\pm 1.67}$ $31.35{\pm 2.32}$- $60.19{\pm 3.86}$
REVS (ours) $80.65{\pm 2.41}$ $97.22{\pm 1.04}$- $\underline{68.98{\pm 3.6}}$
Table 3: Unlearning effectiveness and model integrity results in GPT-J-6B. REVS is superior in almost all cases.
Method Resistance Score $\uparrow$ Logit Lens@100 $\uparrow$ Delta@100 $\uparrow$ Perturb@100 $\uparrow$
SSN Unedited $0{\pm 0.0}$ $0{\pm 0.0}$ $95.12{\pm 0.82}$ $26.5{\pm 5.26}$
FT-L $78.25{\pm 3.79}$ $71.03{\pm 7.07}$ $97.17{\pm 0.74}$ $72.27{\pm 4.82}$
MEMIT $\underline{93.52{\pm 1.76}}$ $\underline{97.48{\pm 2.01}}$ $\underline{97.88{\pm 0.64}}$ $\underline{90.93{\pm 3.53}}$
Max-Entropy $43.70{\pm 10.35}$ $32.90{\pm 9.61}$ $96.77{\pm 0.91}$ $37.23{\pm 9.47}$
Head-Projection $84.55{\pm 2.17}$ $80.22{\pm 2.28}$ $96.20{\pm 0.49}$ $79.42{\pm 4.57}$
RMU $35.60{\pm 8.34}$ $19.90{\pm 6.08}$ $96.45{\pm 0.79}$ $45.58{\pm 9.72}$
NPO-KL $63.17{\pm 7.83}$ $52.58{\pm 10.87}$ $96.80{\pm 0.71}$ $56.23{\pm 7.33}$
REVS (ours) $99.12{\pm 3.56}$ $99.12{\pm 0.51}$ $98.55{\pm 0.2}$ $98.97{\pm 1.46}$
Emails Unedited $0{\pm 0.0}$ $0{\pm 0.0}$ $83.8{\pm 0.67}$ $44.2{\pm 4.11}$
FT-L $39.22{\pm 2.67}$ $20.32{\pm 1.64}$ $84.02{\pm 1.13}$ $65.57{\pm 6.43}$
MEMIT (modified) $\underline{80.73{\pm 1.7}}$ $79.62{\pm 2.31}$ $86.17{\pm 0.39}$ $\underline{77.12{\pm 3.86}}$
Max-Entropy $54.76{\pm 2.27}$ $34.98{\pm 1.94}$ $\bm{87.9{\pm 0.16}}$ $67.7{\pm 41.6}$
Head-Projection $79.92{\pm 1.62}$ $\bm{82.55{\pm 2.63}}$ $81.80{\pm 0.77}$ $75.88{\pm 2.81}$
RMU $29.12{\pm 18.87}$ $15.93{\pm 13.72}$ $84.53{\pm 0.97}$ $62.25{\pm 8.64}$
NPO-KL $44.43{\pm 3.84}$ $26.28{\pm 3.17}$ $82.63{\pm 0.90}$ $58.23{\pm 5.14}$
REVS (ours) $83.48{\pm 1.14}$ $\underline{81.05{\pm 1.17}}$ $\underline{87.08{\pm 0.25}}$ $82.63{\pm 2.63}$
Table 4: Average results for extraction resistance in GPT-J-6B. REVS is more robust to extraction attacks.
REVS outperforms all baselines on both the Unlearning Score and the Resistance Score for the GPT-J-6B model across both the SSN and Emails datasets (Tables 3 and 4). Across individual metrics, REVS consistently achieves the best or second-best performance.
Efficacy
On the SSN dataset, both REVS and MEMIT demonstrate strong unlearning effectiveness, significantly outperforming other baselines. On the Email dataset, while both REVS and Head-Projection show superior effectiveness, REVS achieves near-perfect efficacy.
Generalization
REVS strongly outperforms competing methods in generalization capability, demonstrating robust unlearning across varied contexts even when unlearned on single prompt-target pairs.
Specificity.
On both the SSN and Emails datasets, REVS achieves the second-best score on specificity while maintaining the overall highest Unlearning Score.
General capabilities
Even after few-shot learning, the baseline unedited model performed near random chance on MMLU, achieving scores of $27$ for Emails and $26$ for SSN, where $25$ represents random performance. Consequently, we did not evaluate MMLU and GSM8K scores for GPT-J-6B.
Extraction resistance
As shown in Figures 5, LABEL: and 4, REVS achieves superior Resistance Score, consistently outperforming all baselines including MEMIT and Head-Projection across both datasets. On the SSN dataset, REVS significantly outperforms all other baselines with near-perfect scores across all attacks. On the Emails dataset, REVS achieves best or second-best performance with comparable scores on individual attacks.
A.4 Analysis
Impact of candidate token size
As shown in Figures 6 and 6, REVS maintains superior performance compared to MEMIT across varying candidate token set sizes ($|C|$).
Efficacy/Specificity trade-off
Figures 6 and 6 illustrate the trade-off between efficacy and specificity. REVS can achieve near-perfect specificity at low efficacy levels across both datasets, while MEMIT’s specificity saturates around $0.65$ for the Emails dataset. However, MEMIT shows more consistent specificity scores across configurations, while REVS exhibits larger efficacy-specificity trade-offs.
Scaling with number of edits
As shown in Figures 6 and 6, both methods show declining Unlearning Scores as edit counts increase. REVS achieves near-perfect scores with few edits on the Emails dataset, while MEMIT saturates earlier. On the SSN dataset, both methods show high variance with few targets, but MEMIT demonstrates better stability with increasing targets. Higher scores are achievable by optimizing hyperparameters for larger edit counts, as current parameters were optimized for 100 targets.
Figure 5: Results for Extraction Resistance on GPTJ-6B for $k=100$. REVS is more robust to extraction attacks. Refer to Table 4 for the detailed results.
Figure 6: GPT-J-6B; Top row: SSN dataset; Bottom row: Email dataset. From left to right: (a+d) efficacy vs candidates size, (b+e) efficacy vs. specificity trade-off, (c+f) Unlearning Score vs. number of targets for Emails (c) and number of targets for SSN (f). Mean results with confidence intervals.
Appendix B REVS
B.1 Implementation and Experimental Details
In this section, we provide further implementation details for REVS, describing the hyper-parameter search and experiments conducted in more detail.
We perform an extensive hyper-parameter search was performed, focusing on the maximum number of neurons, neuron selection strategy, desired rank of the target token $r_{n}$ in the neuron $\vec{n}$, and desired rank of the target token $r_{d}$ in the hidden state of each layer $\vec{h}=FF_{2}\vec{a}$.
For the Llama-3-8B model, the hyperparameters are: A maximum of $n_{max}=130$ neurons (SSN) and $n_{max}=45$ neurons (Email) were edited per layer. The neurons were chosen by first selecting the top $k=1000$ neurons with the highest activation values. The desired rank $r_{h}$ within the residual hidden layer was set to $r_{h}=750$ with an error margin of $\epsilon_{r_{h}}=250$. The desired rank $r_{n}$ within the edited neurons was set to $r_{n}=105000$, with an error margin $\epsilon_{r_{n}}=5000$, where the maximum possible rank for any token in the Llama-3-8B vocabulary is $128,256$.
For GPT-J-6B model and SSN and Email datasets, the best configuration found through this tuning process was as follows: A maximum of $N_{max}=30$ neurons were edited per layer. The neurons were chosen by first selecting the top $k=100$ neurons with the highest activation values, and then from within these neurons, the ones with the highest rank for the target token when projected to the vocabulary logit space were selected. For the target token, the desired rank $r_{h}$ within the residual hidden layer was set to $r_{h}=2600$ with an error margin of $\epsilon_{r_{h}}=1400$. The desired rank $r_{n}$ within the edited neurons was set to $r_{n}=37500$, with an error margin $\epsilon_{r_{n}}=7500$, where the maximum possible rank for any token in the GPT-J-6B vocabulary is $50400$. In our hyperparameter sweep, we considered the following ranges: the number of neurons from 10 to 1000 and the activation filter size from 10 to 2000. Additionally, we explored alternative strategies for token selection (Section E.2) and neuron selection (Section D.2).
Original Text Unlearned with REVS
[…]
.c \/ Copyright (C) 1995-1998 Eric Young (e***@cryptsoft.com) * All rights reserved. .c \/ Copyright (C) 1995-1998 Eric Young ([email protected]) * All rights reserved.
[…]
Table 5: Randomly selected examples showing relevant parts of sentences containing naturally memorized email addresses by Llama-3-8B. Original email addresses are censored for privacy reasons. After unlearning with REVS, email addresses are either replaced with plausible alternatives or with "[email protected]" while maintaining natural and coherent text that preserves the original context.
Retain Email Addresses After Unlearning with REVS
The official customer support email address of Microsoft is support@microsoft.com.
If your question is not answered on the Intercountry Adoption website, reach out to the Office of Children’s Issues by email at adoption@state.gov or by phone at 1-888-407-4747.
If you get an email "from the IRS" that you didn’t consent to receive, don’t click on any links. Forward the email to phishing@irs.gov and then delete it.
To email the president in the White House, send an email to president@whitehouse.gov.
Table 6: Demonstration of model’s ability to retain and generate public email addresses after unlearning target-specific private email addresses, showcasing the selective unlearning capability of REVS.
B.2 Editing Neuron
The EditNeuron step, as detailed in Algorithm 2, aims to adjust the logit value of the target token $t$ such that it achieves the desired rank $r_{n}$ in the vector of logits $\vec{v}$. This is accomplished through an iterative process that adjusts the logit value for $t$ and projects the updated logits back and forth between the vocabulary space and the hidden state space until the target rank is achieved.
The iterative approach is necessary because the unembedding matrix $U$ is not invertible. Instead, we use its pseudoinverse, $U^{\dagger}$, which introduces approximation errors. Consequently, when updating the logit value of $t$ to reflect the desired rank $r_{n}$, the projection of the modified logits back to the neuron space does not guarantee that the edited neurons $\vec{n}^{*}$ will yield the correct rank for $t$. This necessitates an iterative process to converge on a logit value $l_{t}$ that achieves the desired rank within a specified tolerance.
The process begins by initializing the logit value of $t$ to $-10$, corresponding to a low probability, and therefore low rank in $\vec{v}$. At each iteration, the logit value is adjusted by scaling it with either an increasing factor $\delta_{h}$ or a decreasing factor $\delta_{l}$, depending on whether the current rank of $t$ is lower or higher than the desired rank $r_{n}$. These factors control the adjustment rate to ensure efficient convergence. After each adjustment, the updated logits are projected back to the hidden state space and then to the vocabulary space to reevaluate the rank of $t$. This process continues until the rank of $t$ is within an acceptable margin of error $\epsilon$ relative to $r_{n}$.
1:procedure EditNeuron($\vec{n},t,r_{n},\epsilon$)
2:     $U\leftarrow\text{Unembedding Matrix}$
3:     $U^{\dagger}\leftarrow\text{Pseudo-inverse of }U$
4:     $l_{t}\leftarrow-10$triangleright Initial low logit value for target token
5:     repeat
6:         $\vec{v}\leftarrow U\vec{n}$triangleright Project neuron to logit space
7:         $\vec{v}_{t}\leftarrow l_{t}$triangleright Update logit value of $t$ in $\vec{v}$
8:         $\vec{n}\leftarrow U^{\dagger}\vec{v}$triangleright Project updated logits to neuron
9:         $\vec{v}\leftarrow U\vec{n}$triangleright Project updated neuron to logits
10:         $r_{t}\leftarrow\mathrm{Rank}(t,\vec{v})$triangleright Get rank of $t$ in projected logits
11:         if $|r-r_{n}|\leq\epsilon$ then
12:              break $\triangleright$ Stop if rank is within $\epsilon$ of desired rank
13:         else if $r_{t}>r_{n}$ then
14:              $l_{t}\leftarrow l_{t}*1.3$triangleright Decrease $l_{t}$ if rank is too high
15:         else
16:              $l_{t}\leftarrow l_{t}*0.8$triangleright Increase $l_{t}$ if rank is too low
17:         end if
18:     until true
19:     return $\vec{n}$triangleright Return updated neuron value
20:end procedure
Algorithm 2 Edit Neuron Algorithm
B.3 Qualitative Unlearning Examples
We demonstrate the Unlearning Effectiveness and Model Integrity of REVS through qualitative examples in Tables 5 and 6. Table 5 presents randomly sampled instances from our dataset, showcasing how the unlearning process handles private email addresses. After applying our method, the model generates plausible alternatives to the original private email addresses. Specifically, we observe two primary transformation cases: (1) replacing the original email address with a similar but different email address that maintains the contextual integrity of the original text; (2) replacing the specific email with a generic [email protected] placeholder while preserving the overall coherence and meaning of the surrounding text.
To complement our demonstration, Table 6 shows select public email addresses, exemplifies how the model maintains generation capabilities for publicly available contact information after the unlearning process.
Appendix C Baselines Implementation and Experimental Details
We performed an extensive hyper-parameter search for each baseline on each dataset. For all baselines, we conducted a sweep over a wide range of key hyper-parameter values, ensuring comprehensive exploration of the parameter space. The specific hyperparameter values are detailed below for each model and dataset.
C.1 MEMIT Baseline
We used a modified version of MEMIT (Meng et al., 2023), originally proposed for inserting new knowledge into language models by editing the $FF_{2}$ matrices in model MLPs. In our case, we optimized the objective to decrease the probability of generating the target tokens instead of increasing it, to facilitate unlearning. The original MEMIT method was proposed for editing memorized facts of the form (subject s, relation r, object o), e.g., (s = Michael Jordan, r = plays sport, o = basketball), where the subject is used for calculating the updated weights. However, the targeted sensitive information might not have an explicit subject (e.g., "Send me an email with your decision soon as you get this; david.lewis@email.com"), which is the case for many samples in the Email dataset (Table 11). Furthermore, the targeted sensitive information may be memorized by various inherently different prompts where each sentence might have unique subject (e.g., "David’s email is: david.lewis@email.com" and "Contact Lewis at: david.lewis@email.com). To address this challenge, the following tuple was used as the input to the modified model: (s = last characters $n_{max\_len}$ of the prompt, r = <empty>, o = target token to unlearn).
For the Llama-3-8B model, we applied the method with specific hyperparameters: on the SSN dataset, we focused on layer $3$ with a loss break of $1\times 10^{-5}$, loss prediction probability coefficient of $100$, learning rate of $0.01$, and $5$ gradient steps with a maximum prompt length of $n_{max\_len}=20$. For the EMAIL dataset, we maintained similar settings, adjusting only the maximum prompt length to $n_{max\_len}=100$. For the GPT-J-6B model, the following hyperparameters were used: on the SSN dataset, we targeted layers $3$-$8$ with a loss break of $0.1$, loss prediction probability coefficient of $20$, learning rate of $0.05$, $25$ gradient steps, and a maximum prompt length of $n_{max\_len}=200$. The EMAIL dataset configuration was similar, with a reduced maximum prompt length of $n_{max\_len}=50$. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
C.2 FT-L Baseline
Constrained Fine-tuning (FT-L) (Zhu et al., 2020) fine-tunes the $FF_{2}$ using gradient ascent to minimize the probability of generating the target tokens $t$. FT-L imposes an $L_{\infty}$ norm constraint to limit the overall impact on the model’s knowledge. For the FT-L baseline, we employed different hyperparameter configurations across models and datasets. For both models we targeted all layers. On Llama-3-8B, For the SSN dataset, we used a loss break of $8.362\times 10^{-2}$, learning rate of $1.12\times 10^{-7}$, norm constraint of $3.3\times 10^{-5}$, and $10$ gradient steps. The EMAIL dataset configuration used a loss break of $1.225\times 10^{-3}$, learning rate of $1.75\times 10^{-7}$, norm constraint of $4.17\times 10^{-5}$, and the same $10$ gradient steps. On GPT-J-6B model, for the SSN dataset, we applied a loss break of $9.98\times 10^{-2}$, learning rate of $3.78\times 10^{-7}$, norm constraint of $7.1\times 10^{-5}$, and $10$ gradient steps across layers 00-$27$. The EMAIL dataset configuration adjusted these values, using a loss break of $1.49\times 10^{-3}$, learning rate of $1.12\times 10^{-7}$, norm constraint of $3.84\times 10^{-5}$, while maintaining $10$ gradient steps. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $1\times 10^{-8}$ to $5\times 10^{-3}$, loss break from $1\times 10^{-3}$ to $1\times 10^{-1}$, and norm constraint from $1\times 10^{-6}$ to $1\times 10^{-2}$, and number of gradient steps from 5 to 25.
C.3 NPO-KL Baseline
NPO-KL (Zhang et al., 2024) implements Negative Preference Optimization by maximizing the prediction probability on the target while minimizing the prediction probability of alternatives. This version adds a KL divergence loss between the model’s outputs before and after unlearning to minimize interference. We modified the NPO implementation to target a single sensitive token instead of the entire text. This was done to unlearn only the relevant sensitive target token and avoid unnecessary changes to the model’s weights. We used the KL version of the baseline due to its potential to preserve more information within the model that we do not target to unlearn. For the Llama-3-8B model, On the SSN dataset, we used a beta of $100$, KL coefficient of $1\times 10^{-2}$, learning rate of $5.8\times 10^{-7}$, and $2$ epochs with a maximum prompt length of $n_{max_{l}en}=100$. The EMAIL dataset configuration differed, maintaining a beta of $100$ but increasing the KL coefficient to $100$, with a learning rate of $6.3\times 10^{-7}$ and $3$ epochs. On the GPT-J-6B, for the SSN dataset, we employed a beta of $100$, KL coefficient of $1$, learning rate of $1.32\times 10^{-6}$, $2$ epochs, and a maximum prompt length of $n_{max_{l}en}=100$. The EMAIL dataset configuration used a beta of $0.1$, KL coefficient of $1\times 10^{3}$, learning rate of $6.1\times 10^{-7}$, and $3$ epochs. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-7}$ to $1\times 10^{-3}$, number of epochs from 1 to 5, and values for regularization parameters $\beta$ and KL coefficient of $[0.01,0.1,1,10,100,1000]$.
C.4 RMU Baseline
RMU (Li et al., 2024) (Representation Misdirection for Unlearning) perturbs model activations of a selected layer on the target while preserving model activations on benign data. Similar to NPO-KL, we modified RMU to target a single sensitive token. On Llama-3-8B, for SSN, we used an alpha of $1$, layer ID $7$, layers $5$-$6$-$7$, learning rate of $6.819\times 10^{-4}$, $2$ epochs, and a steering coefficient of $1$. The EMAIL dataset configuration shifted to an alpha of $1000$, maintaining the same layer and learning rate specifications, but increasing the steering coefficient to $1000$. On GPT-J-6B model, for the SSN dataset, we applied an alpha of $100$, layer ID $7$, layers $5$-$6$-$7$, learning rate of $5.19\times 10^{-4}$, $2$ epochs, and a steering coefficient of $1000$. The EMAIL dataset configuration used an alpha of $329$, the same layer specifications, a learning rate of $7.74\times 10^{-4}$, $2$ epochs, and a reduced steering coefficient of $1.61\times 10^{-1}$. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $1\times 10^{-5}$ to $5\times 10^{-2}$, number of epochs from 1 to 5, and values for regularization parameters $\alpha$ and steering coefficient of $[1,10,100,1000]$.
C.5 Head-Projection Baseline
We followed the original implementation with a modified objective to prevent the target from appearing in the top $k=100$ logit distribution elements. For the Llama-3-8B model, we targeted layers $3$-$5$ and Patil layers $15$-$31$. The SSN dataset configuration used a loss break of $1.9\times 10^{-3}$, loss prediction probability coefficient of $3.08\times 10^{-2}$, learning rate of $1.03\times 10^{-2}$, and $10$ gradient steps. The EMAIL used a loss break of $7.6\times 10^{-4}$, loss prediction probability coefficient of $2.58\times 10^{-1}$, learning rate of $9.19\times 10^{-3}$, and $25$ gradient steps. For the GPT-J-6B model we targeted layers $3$-$8$ and Patil layers $15$-$27$. For the SSN dataset we used a loss break of $6.596\times 10^{-2}$, loss prediction probability coefficient of $3.271$, learning rate of $2.045\times 10^{-1}$, and $25$ gradient steps. The EMAIL dataset configuration used similar layer targeting but modified the hyperparameters, with a loss break of $6\times 10^{-4}$, loss prediction probability coefficient of $6.733\times 10^{-1}$, learning rate of $1.868\times 10^{-1}$, and $5$ gradient steps. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
C.6 Max-Entropy Baseline
For the Llama-3-8B model, we targeted layers $3$-$5$ and Patil layers $15$-$31$. The SSN dataset configuration employed a loss break of $1.99\times 10^{-2}$, loss prediction probability coefficient of $20.5$, learning rate of $7.6\times 10^{-3}$, and $25$ gradient steps. The EMAIL dataset adjusted these parameters, using a loss break of $1.75\times 10^{-3}$, loss prediction probability coefficient of $1.35\times 10^{-2}$, learning rate of $9.2\times 10^{-3}$, and $5$ gradient steps. For the GPT-J-6B model, we targeted layers $3$-$8$ and Patil layers $15$-$27$. For the SSN dataset, we used a loss break of $7.0559\times 10^{-2}$, loss prediction probability coefficient of $25.9$, learning rate of $1.937\times 10^{-1}$, and $5$ gradient steps. The EMAIL dataset used a loss break of $3.634\times 10^{-2}$, loss prediction probability coefficient of $1.907\times 10^{-2}$, learning rate of $6.65\times 10^{-2}$, and $10$ gradient steps. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
Appendix D Ablations
D.1 Alternative Neuron Selection Strategies
We explored several alternative methods for neuron selection, besides our hybrid approach that considers both neuron activation and association with the target token (Section 3.1.2). These alternatives include: (a) selecting the most activated neurons given the prompt, (b) selecting neurons solely based on their association with the target token, (c) neurons with highest gradients with respect to the prompt (Dai et al., 2022), and (d) random neuron selection as a baseline. As shown in Table 7, our evaluation on the Emails dataset on GPT-J-6B, using the same seed (00) that was used for finding the best hyper-parameters, demonstrated that the hybrid approach (rank & activation) outperformed these alternatives in achieving superior results in both metrics of unlearning effectiveness and model integrity.
Neuron Selection Method Unlearning Score Resistance Score
Random 00 00
Gradient-based $71.1$ $78.2$
Rank $22.4$ $29.2$
Activations $68.8$ $81.8$
Rank & Activations (hybrid) $\bm{75.2}$ $\bm{84.2}$
Table 7: Ablation experiments on REVS for neuron selecting method. Our hybrid approach yields the best Unlearning Effectiveness and Attacks Resistance with Model Integrity. GPT-J-6B on Email dataset.
D.2 Alternative Neuron Editing Strategies
We explore an alternative neuron editing strategy inspired by DEPN, where the target neurons are directly zeroed out instead of applying our proposed rank-based editing approach. As shown in Table 8, zeroing neurons results in significantly inferior performance compared to REVS.
The findings indicate that directly zeroing the selected neurons fails to achieve the desired unlearning objectives, underscoring the need for a more nuanced and precise approach to model editing. The superior performance of our method highlights the importance of employing a targeted and surgical strategy for neuron editing.
Method Unlearning Score $\uparrow$ Efficacy@100 $\uparrow$ Specificity $\uparrow$ Resistance Score $\uparrow$
$ZERO_{n=5}$ 00 $73.8$ 00 $79.2$
$ZERO_{n=2}$ $14.3$ $26.3$ $9.7$ $47.5$
$ZERO_{n=1}$ $11.2$ $6.8$ $31.9$ $17.3$
REVS (ours) $\bm{83.5}$ $\bm{81.1}$ $\bm{87.1}$ $\bm{82.6}$
Table 8: Ablation experiments - Zeroing target neurons (vs. rank editing): Substantial performance superiority of REVS over baseline approaches, with $n$ indicating number of zeroed neurons. GPT-J-6B on Email dataset
Appendix E Token Selection
E.1 Details and Examples
All methods in this work unlearn the same target tokens from sequences containing sensitive information. Our approach strategically selects the rarest tokens $T\subseteq S$ within each sequence $S$ to minimize unnecessary model perturbation while effectively isolating sensitive content. Token rarity is approximated using token ID assignments, where higher ID values indicate rarer tokens.
In our experiments, we unlearn two tokens ($|T|=2$) per target sequence. For example, consider the email address lewis.david@email.com, tokenized as [le, wis, .d, avid, @email, .com] with token IDs [273, 49143, 962, 15567, 72876, 916]. Here, we select wis and avid as the rarest tokens while excluding common domain-related tokens like @email.com.
For the URL dataset, we exclude common substrings such as [http, https, www, ://, /, ", -] from unlearning, as they do not encapsulate meaningful information and remain easily inferable. Similarly, in the SSN dataset, we exclude [-] for the same reason.
This strategy ensures the removal of the most unique, information-dense tokens that are most likely to contain sensitive identifiers while minimizing unnecessary model perturbation.
E.2 Alternative Token Selection Strategies
We investigated token selection strategies for unlearning on the GPT-J-6B Email dataset. Table 9 shows that selecting the rarest tokens yields the best performance across all metrics. The rarest token approach significantly outperforms other strategies like most frequent, first, or random token selection, demonstrating substantially higher effectiveness in unlearning.
Token Selection Unlearning Score $\uparrow$ Efficacy@100 $\uparrow$ Specificity $\uparrow$ Resistance Score $\uparrow$
Most Frequent 55.6 67.7 47.22 70.6
First 65.6 86.8 52.77 77.2
Random 72 89 60.4 80.5
Rarest 75.2 96.1 61.8 83.9
Table 9: Ablation experments on REVS over tokens selection method. Selecting the rarest tokens yields the best Unlearning Effectiveness and Attacks Resistance with Model Integrity. GPT-J-6B on Email dataset.
Appendix F Dataset Curation
Existing work on removing sensitive information (Patil et al., 2024) lacks datasets for evaluating unlearning methods on actual sensitive information, particularly on organically memorized sensitive information.
While datasets such as Who’s Harry Potter (Eldan and Russinovich, 2023), WMDP (Li et al., 2024), and TOFU (Maini et al., 2024) are widely used, they focus on concept unlearning. Concept unlearning differs fundamentally from the challenge of unlearning sensitive information, where the goal is to obfuscate a single piece of sensitive information rather than broad knowledge of a concept, rendering these datasets unsuitable for our specific goals.
To the best of our knowledge, no publicly available dataset contains sensitive information that models naturally memorize. To address this gap, we curated a novel English language dataset specifically designed to evaluate unlearning methods on real-world, naturally memorized for each model. To provide a more comprehensive and diverse evaluation, we also curated a synthetic English language dataset in which the sensitive token sequences are numbers (as opposed to the character-based sequences in the email and URL datasets).
F.1 Naturally Memorized Dataset
Curating a dataset of naturally memorized sensitive information involves identifying sensitive instances in the model’s training data and verifying their memorization. However, determining both the sensitivity of an instance and its memorization is computationally expensive and requires specialized tools like Presidio (Microsoft, 2024).
Existing work, such as Elazar et al. , use regex-based filtering on datasets like The Pile (Gao et al., 2020) to identify sensitive information. However, these methods produce many false positives—instances flagged as sensitive but that are not genuinely so.
To address this, we leveraged the “Training Data Extraction Challenge” [^2] https://github.com/google-research/lm-extraction-benchmark/blob/master/detailed_description.pdf, a subset of The Pile known to contain sensitive information. This subset comprises 15,000 sentences in which GPT Neo exhibited verbatim memorization of sentence suffixes given their prefixes. We used this subset as a strong candidate for naturally memorized sensitive data.
We analyzed the dataset using Presidio (Microsoft, 2024) to identify various types of personally identifiable information (PII). For each identified PII instance, we tested whether the model could reproduce it exactly when given the original context prefix. The distribution of identified and memorized PII types by Llama 3 8B is summarized in Table 10.
PII Type Identified Memorized
URLs 17,233 203
Email addresses 1,631 205
Phone numbers 357 0
IP addresses 55 0
Table 10: Distribution of identified and memorized PII types by Llama 3 8B in the dataset
F.1.1 Emails Dataset
To curate the email dataset, we used Presidio to identify sentences containing email addresses. We then test memorization by providing the model with the text preceding each email address and checking whether it could generate the exact email within the next 50 tokens using greedy decoding. The target token sequence for unlearning is defined as the substring preceding “@email.com.”
Our results show that Llama 3 8B memorized 205 email addresses from the dataset, while GPT J 6B memorized 288. Sample sentences from this dataset, with emails obfuscated, are presented in Table 11.
F.1.2 URLs Dataset
For the URL dataset, we identified sentences containing URLs using Presidio. We filtered out URLs shorter than 40 characters to exclude generic addresses (e.g., “https://www.gnu.org”) and URLs containing the word “License.” Our analysis found that Llama 3 8B memorized 203 URLs meeting these criteria. We did not evaluate URL memorization for GPT J 6B.
Handling of sensitive data
While The Pile is a publicly available dataset, appropriate measures were taken to handle the sensitive private information present in the dataset, which contained real email addresses. The raw dataset was accessed only by the authors. At no point were any individual email addresses or other PII exposed or included in the research outputs or publicly released datasets. We will disseminate the email addresses memorized by GPT-J upon request for research purposes if researchers state they will adhere by the same standards.
Sentence
1 Copyright 2004-2006, 2009-2010 Colin Green (sharp***@gmail.com)
2 Let Gus know by sending an email to g***@flyin***.com.
3 If you need assistance with this program, you may contact: m***@laptop.org.
4 You can contact SugarCRM, Inc. headquarters at c***@sugarcrm.com.
5 Authors: # Mike Auty <mike.***@gmail.com>
6 Please send bug reports and support requests to l***@saillard.org.
Table 11: Example sentences containing obscured email addresses from the dataset.
F.2 Synthetic SSN Dataset
The Synthetic SSN dataset was generated by prompting Claude 3 Sonnet (Anthropic, 2024) to create template sentences on various topics that could contain SSNs (employment records, tax documents, finance, government records, and medical records), with placeholders for names, dates, and SSNs. We then randomly assigned fake sensitive information to each of the placeholders. Each base model is fine-tuned to memorize all instances. When constructing the dataset, we assign 20 unique SSN targets across $100$ templates, such that each target SSN has $5$ different prefixes. This allows evaluating generalization: we unlearn using only one prompt that generates the target token sequence, and evaluate on four additional prompts that also generate it. The remaining $100$ sentences were used to evaluate the method’s specificity, that is, whether it affects memorization of SSN targets that were not unlearned. In this dataset, the target tokens to unlearn were chosen only from numeric tokens, as the SSN structure contains ’-’ (e.g., 123-45-6789).
We curated the synthetic SSN dataset by using the following prompt with Claude 3 Sonnet:
You are a synthetic data generation tool. I need you to create template sentences across different domains that could potentially contain sensitive personal information such as names, social security numbers, and dates. Instead of using real data, use placeholder tags like [NAME], [SSN], and [DATE] within sentences. For each domain listed below, provide 20 full sentences written naturally, as if they could appear in real documents from that domain. Ensure the placeholders [NAME] and [SSN] appear mid-sentence or at the end, not at the beginning. Use diverse sentence structures and phrasings for variety, and do not include any real personal data, only placeholder tags.
Example sentences from the medical domain: Adjusting the medication dosage on [DATE] to better manage a thyroid condition, the endocrinologist treated [NAME], bearer of SSN [SSN]. Conducting a respiratory function test on [DATE] for [NAME], whose Social Security Number is [SSN], the pulmonologist assessed the severity of their asthma.
The domains for which diverse template sentences are required include: Employment Records, Tax Documents, Financial Records, Government Records, and Medical Records.
Please proceed by providing 20 unique and diverse sentences per domain in the specified format using the placeholder tags. Let me know if you need any clarification on the instructions.
We kept asking the model to generate examples until we had a total of 200 sentences.
We then randomly assigned fake sensitive information to each placeholder. To populate the placeholders, we used Claude 3 Sonnet to generate a list of $40$ unique names (since each sentences will be assigned with the same name and SSN for allowing evaluating generalization). Additionally, we generated $200$ random dates and $40$ random Social Security Numbers (SSNs). To get the model to memorize the sentences containing the sensitive information, we fine-tuned the model with a low learning rate of $62e-$ 6 to minimize unintended impact. The fine-tuning process ran for 4 epochs with a batch size of $1$. Table 12 shows 10 example sentences from this dataset.
Sentence
1 During a routine check-up on 23-3-2021, Michael Choi, with the Social Security Number 135-54-3368, discussed their recent test results and any necessary follow-up steps with the attending physician.
2 On 2-2-2021, Michael Choi, whose Social Security Number is 135-54-3368, participated in a group therapy session focused on coping mechanisms for managing chronic pain.
3 At a scheduled appointment on 19-6-2016, Michael Choi, with an SSN of 135-54-3368, received guidance from a physical therapist on exercises to improve mobility and strengthen core muscles.
4 During a visit on 22-10-2018, Michael Choi, whose Social Security Number is 135-54-3368, consulted with a dermatologist about a concerning mole and underwent a biopsy procedure.
5 On 22-10-2019, Michael Choi, with the Social Security Number 135-54-3368, attended a seminar on the importance of regular preventive screenings and early detection of various health conditions.
6 The patient, a 45-year-old Mark Evans, with the Social Security Number 404-61-5777, visited the clinic on 28-11-2015 for a routine physical examination.
7 During a follow-up appointment on 8-1-2014, the physician reviewed the recent test results with Mark Evans, whose Social Security Number is 404-61-5777.
8 On 12-12-2014, the physical therapist designed a customized exercise plan for Mark Evans, with an SSN of 404-61-5777, to aid in their recovery from a knee injury.
9 The nutritionist provided dietary recommendations on 7-5-2021 to Mark Evans, whose Social Security Number is 404-61-5777, to help manage their high cholesterol levels.
10 At the mental health clinic on 13-6-2013, Mark Evans, with the Social Security Number 404-61-5777, participated in a group therapy session focused on stress management techniques.
Table 12: Examples from our synthetic dataset containing SSN. Each group of 5 sentences shares a unique SSN. Unlearning efficacy is measured on the SSN given the targeted prompt, while generalization is evaluated as efficacy using the remaining 4 unseen sentences that contain the same SSN.
Appendix G Perturbation Attack (PA)
The Perturbation Attack (PA) is a white-box attack that aims to fool the language model by introducing subtle perturbations to the input prompts, which can be adapted to a black-box setting by using only the last hidden state. In this attack, the original prompts are modified by randomly inserting characters at various positions. We experimented with different perturbation strategies, including varying the types of characters inserted and the number of insertion points. Our findings indicate that the most effective approach is to insert spaces into the original prompt at $10$ different random indices, as well as inserting a space immediately after the prompt. Similar to the Logit-Lens Attack (LLA), the candidate set $C_{\textrm{PA}}$ for the PA is obtained by projecting each layer’s residual hidden state to the vocabulary space and considering the top-$k$ highest and lowest $k$ ranked tokens as candidates.
Appendix H Hardware Details
The experiments described in this work were conducted on a computing system equipped with 32 Intel(R) Xeon(R) Gold 6430 CPUs operating at 1.0TB RAM. The underlying hardware consisted of NVIDIA RTX 6000 Ada Generation GPUs, each equipped with 49GB of VRAM. For running methods on the Llama-3-8B model, we used the following GPU configurations for distinct methods: REVS: 1 GPU, MEMIT: 2 GPUs, FT-L: 4 GPUs, RMU: 3 GPUs, Head Projection: 4 GPUs, Max-Entropy: 4 GPUs, and NPO-KL: 5 GPUs. Notably, while some methods required multiple GPUs for efficient processing, REVS maintained computational efficiency with minimal GPU resources, since it does not need to compute the gradient for the unlearning. Runtime on the Email Dataset with Llama-3-8B for each method is as follows: REVS takes 39 minutes, FT-L completes in 8 minutes, MEMIT, Max-Entropy and Head-Projection each take 13 minutes. Meanwhile, RMU and NPO-KL both have a runtime of 9 minutes.